OpenStudy (anonymous):

Graph has a vertex at (−1,−2) and passes the point (0,−1). Find the quadratic function.

OpenStudy (mertsj):

\[y=a(x-h)^2+k\]

OpenStudy (mertsj):

Plug in and solve for a. h = -1, k = -2, x = 0, y = -1

OpenStudy (anonymous):

hmm i got -3(x+1)^2 + 3

OpenStudy (anonymous):

a(x+1)^2 + 4 a(0+1)^2+4 a(1)+4 -1=1a+4 -5==1a a=-5

OpenStudy (anonymous):

-5(x+1)^2 + 4??

OpenStudy (mertsj):

-1=a(0+1)^2-1 -1=a(1)-1 a=